OpenStudy (linda):

f(x)=2cos(x)in the interval 0 is less than or equal to x less than or equal to 2 pi. if g(x) is a shift of f(x) a distance of pi over 2 to the right, write an equation for g(x). anyone know????

OpenStudy (anonymous):

All you would want to do is shift the x on the inside. So g(x)=2cos(x-2). So this takes whatever x value you have, removes two from it, and moves it two over from where it would normally be (f(x)).

OpenStudy (anonymous):

2cos(x-pi/2)

OpenStudy (anonymous):

yes, it will be \[f(x)=2\cos(x-\frac{\pi}{2})\]better known as \[sin(x)\]

OpenStudy (anonymous):

Sorry, I didn't see the pi/2 D: But yes, g(x)=2cos(x-(pi/2))

OpenStudy (anonymous):

i should say better known as \[2\sin(x)\]

OpenStudy (anonymous):

both maleviolence and sleek-feathered one (?!) are right, but you should convince yourself that \[\cos(x-\frac{\pi}{2})=\sin(x)\]

OpenStudy (linda):

thank you everyone! and then the next one is to write an equation involving no shift. would that be 2sin(x)?

OpenStudy (anonymous):

confused. if it is no shift then wouldn't it be the original function?

OpenStudy (linda):

yeah im confused to. idk i think it would be a 2sin(x) but im not sure

OpenStudy (anonymous):

what does the question say exactly?

OpenStudy (anonymous):

sin(x) and cos(x) are complimentary functions which means if you subtract 90 degrees (the compliment) you get the other function. so cos(x-(pi/2))=sin(x)

OpenStudy (anonymous):

To see this, graph sin(x) and cos(x) and notice that one is shifted pi/2 from the other.

OpenStudy (linda):

ohh ok gotcha, thank you

OpenStudy (anonymous):

No problem :P

OpenStudy (anonymous):

you can also show this using the "addition angle" formula if you got there yet